# Bi-VLA: Vision-Language-Action Model-Based System for Bimanual Robotic Dexterous Manipulations

### v1

@@ V Results and Evaluation
We evaluated the performance of the components of our proposed system both separately and together. This approach allowed us to identify the strengths and weaknesses of each component, as well as to understand how the different parts of the system interact and contribute to the overall system performance. We define a task as completed successfully if the generated code by the LLM-based Code Generator is correct before execution by the Motion Controller and if the actions are successfully executed.
In the first part of the experiment, we generated 100 different requests with GPT-3.5 for the system to prepare a specific type of salad. $1/3$ of the requests were for making vegetable salads, $1/3$ for a Russian salad, and the last $1/3$ for a fruit salad. Our proposed system successfully generated the correct executable code for all the requests, achieving a success rate of 100%.
In the second part of the experiment, 100 pictures with different ingredients were captured to evaluate the VLM system, and $1/3$ of the pictures were used to evaluate each of the salads. Half of the pictures had the right ingredients required for each of the salads, and the other half contained some missing ingredients. To assess the performance of the VLM, we checked if the model returned the correct list of all the ingredients in the image and if each mentioned ingredient in the recipe of the salad was detected correctly (that is, the captions on the ingredients were correct in the output image).
For the images with the correct ingredients, the Vision Module achieved a success rate of 83.4% in giving the correct list of ingredients and a success rate of 96.06% in generating the exact captions on the image.
However, for images with missing ingredients, the model sometimes wrongly assigned the caption of a missing ingredient to another present ingredient on the image, leading to lower performance compared to the previous case, with a success rate of 71.22% in giving the correct list of ingredients and 86.53% in generating the exact captions.
It is important to report that from all the experiments, the success rate of the system depends highly on the performance and precision of the visual component; hence, a wrong answer from the Visual Module directly leads to a failure in satisfying the user’s given request.
@@ VI Conclusion
The Bi-VLA system showcases the integration of vision, language understanding, and robotic action to enable dexterous bimanual manipulations in household tasks. Through a series of experiments, we have demonstrated the system’s ability to accurately interpret human instructions, perceive the visual context of ingredients, and execute the necessary actions to assemble the requested salads. The high success rates achieved by the Language and Vision modules highlight the critical role of these components in the overall system performance.
The findings underscore the importance of continued advancements in visual understanding, as the accuracy of the Vision Module directly impacts the system’s ability to successfully complete the given tasks. Future work will focus on enhancing the robustness and versatility of the Vision Module to handle a wider range of visual variations and uncertainties, ultimately leading to more reliable and adaptable robotic assistants.

### v2

@@ V Results and Evaluation
We assessed the performance of our system’s components both individually and collectively, enabling us to identify their strengths and weaknesses and understand their interactions in contributing to overall performance. We define a task as completed successfully if the generated code by the Code Generator is correct before execution by the Motion Controller and if the actions are successfully executed.
  1. 1.
    Firstly, we evaluated the performance of the Language Module by generating 100 different requests with GPT-3.5 for the system to prepare three types of salad, e.g. “I’m really craving a vegetable salad. Could you make one, please?". Specifically, one-third of the requests were for making vegetable salads, another third were for a Russian salad, and the final third was for a fruit salad. Our proposed Language Module successfully generated the correct executable code for all the requests, achieving a success rate of 100%.
  2. 2.
    In the second part of the experiment, we evaluated the vision module’s performance using 100 pictures of different ingredients, with one-third allocated to each salad type. For each salad type, half the pictures included all the required ingredients (Scenario 1), while the other half were missing some (Scenario 2). We evaluated if the model returned the correct list of all the ingredients in the input image and whether each mentioned ingredient from the salad recipe was detected correctly; that is, the captions on the ingredients and the 2D bounding boxes were accurate in the output image.
    • In Scenario 1, where all ingredients were correct, we achieved a 96.06% success rate in detecting ingredients and an 83.4% success rate in generating correct actions.
    • In Scenario 2, the vision module misidentified present ingredients as missing ones, leading to a detection performance of 86.53% and a 71.22% success rate in generating correct actions.
  3. 3.
    For the Action Generation Module, we ensured reliable translation from 2D pixel to 3D world coordinates as described in Section III-C2, enabling the robot’s gripper to successfully grasp objects of varying orientations and perform cutting operations accurately. Therefore, we assume that the success rate of the module is equivalent to the success rate of generating correct 2D bounding box coordinates.
The table below summarizes the main experiment results:
TABLE I: Experiment Results
Experiments Success rate (%)
LLM Module (all scenario)   100%
VLM Module (scenario 1)   96.06%
VLM Module (scenario 2)   86.53%
Action Generation Module (scenario 1 )   83.4%
Action Generation Module (scenario 2)   71.22%
@@ VI Conclusion
In this work, we explore a new approach to dexterous bimanual manipulation by integrating vision, language understanding, and robotic action generation into a unified system named Bi-VLA. Our experiments demonstrate the system’s capability to accurately interpret human instructions, perceive the visual context of the scene, and execute the required actions. The high success rates of the language and vision modules emphasize their crucial roles in overall system performance. Our findings highlight the need for ongoing advancements in visual understanding, as the accuracy of the vision module significantly affects task completion. Future efforts will focus on enhancing the robustness and versatility of the vision module to better handle a variety of visual variations and uncertainties, ultimately making robotic assistants more reliable and adaptable.
Acknowledgements
Research reported in this publication was financially supported by the RSF grant No. 24-41-02039.